OpenStudy (anonymous):

Can someone please help me with Trig Identities? sec^2x = (1-cos^2x)/(1-sin^2x) + csc^2x - cot^2x

OpenStudy (fibonaccichick666):

Can you elaborate please?

OpenStudy (anonymous):

So far I have: \[\frac{ 1}{ \cos ^{2} x} = \frac{ \sin ^{2}x }{ \cos ^{2}x } + \frac{ 1 }{ \sin^{2}x } - \frac{ \cos^{2}x }{ \sin^{2}x }\] I'm stuck at this part.

OpenStudy (fibonaccichick666):

ok so first don't change the left side at all otherwise the proof is a no go. second sin/cos =what?

OpenStudy (anonymous):

Tangent

OpenStudy (fibonaccichick666):

ok so what properties do you have about tan^2

OpenStudy (anonymous):

like tan^2x + 1 = sec^2x?

OpenStudy (fibonaccichick666):

yup

OpenStudy (fibonaccichick666):

but as a side note, are you sure this even works?

OpenStudy (anonymous):

Oh, I get it now! By this, do you mean the whole equation and thank you very much!

OpenStudy (fibonaccichick666):

oh wait a sec I see an easier way

OpenStudy (fibonaccichick666):

Come back, that way was blah this is easy, just make them all into one fraction

OpenStudy (anonymous):

Oh I see! Thank you again!

OpenStudy (fibonaccichick666):

ok np have a good one